Are the following lines parallel, perpendicular, or neither:
a.) 9x - 12y = 17 and 3x - 4y = 17
b.) y = -3 and x = -3
c.) y = 3x + 4 and y = -3x - 2
How do I solve these. Please explain and show work so I can follow along/understand what I am suppose to be doing.


Answer by cira554(7)

you should know that for lines to be perpendicular the product of their slopes should be -1.
for parallel lines their slopes are equal.
therefore you should first rearrange your equations into linear forms that is in the form y=mx+c where m is the slope.
so in a) 9x-12y=17 and 3x-4y=17 the equations are
-12y=-9x+17 and -4y=-3x+17
y=3/4x-17/12 and y=3/4x-17/4
we can there conclude that they are parallel since the slopes are equal that is 3/4.
b)they are neither since we do not even have the respective slopes.
c)y=3x+4 and y=-3x-2 are also neither since the slopes are not equal
that is 3 and -3,and their product is not equal to -1.